# Near-Optimal Adaptive Transforms for LLM Quantization

### v1

@@ 2 Theoretical Derivation
In this section, we focus on the question of finding optimal transforms that minimize the loss of one block $\ell_{\left(i\right)}$. We omit the subscript $\left(i\right)$ to simplify notation.
@@ 2.1 Problem Setup and General Proof Approach
Probabilistic reformulation. We reformulate the problem from a probabilistic perspective. We split the matrices $\bm{W}=\left[\bm{w}_{1},\dots,\bm{w}_{d_{\mathrm{out}}}\right]$ and $\bm{X}=\left[\bm{x}_{1},\dots,\bm{x}_{d_{\mathrm{batch}}}\right]$ into columns $\bm{w}_{k},\bm{x}_{k}\in\mathbb{R}^{d}$ and treat the columns as i.i.d. samples from $d$-dimensional independent distributions $\bm{w}\sim\mathcal{D}_{\mathrm{w}}$ and $\bm{x}\sim\mathcal{D}_{\mathrm{x}}$, respectively. Then the matrices $\bm{W}^{\prime}$ and $\bm{X}^{\prime}$ satisfy $\bm{W}^{\prime}\bm{W}^{\prime\top}=\mathbb{E}_{\bm{w}}\bm{w}\bm{w}^{\top}$ and $\bm{X}^{\prime}\bm{X}^{\prime\top}=\mathbb{E}_{\bm{x}}\bm{x}\bm{x}^{\top}$, respectively. We can reinterpret the blockwise loss as $\ell=\mathbb{E}_{\bm{w},\bm{x}}\left(q\left(\bm{T}_{\mathrm{W}}\bm{w}\right)^{\top}q\left(\bm{T}_{\mathrm{X}}\bm{x}\right)-\bm{w}^{\top}\bm{x}\right)^{2}$.
Unbiased quantization. We assume that $q$ is a stochastic and unbiased quantizer, i.e., for a vector $\bm{\alpha}\in\mathbb{R}^{d}$, the quantization error $\varepsilon\left(\bm{\alpha}\right)=q\left(\bm{\alpha}\right)-\bm{\alpha}$ is a random vector with $\mathbb{E}_{\varepsilon\left(\bm{\alpha}\right)}\varepsilon\left(\bm{\alpha}\right)=\bm{0}$. We further constrain $\bm{T}_{\mathrm{W}}=\bm{T}_{\mathrm{X}}^{-\top}$ such that $q\left(\bm{T}_{\mathrm{W}}\bm{w}\right)^{\top}q\left(\bm{T}_{\mathrm{X}}\bm{x}\right)$ is unbiased with respect to $\bm{w}^{\top}\bm{x}$. Then, using the first-order approximation, we can split $\ell$ into two non-negative terms:
$=$ $\ \mathbb{E}_{\bm{w},\bm{x},\varepsilon\left(\bm{T}_{\mathrm{W}}\bm{w}\right),\varepsilon\left(\bm{T}_{\mathrm{X}}\bm{x}\right)}\left(\bm{x}^{\top}\bm{T}_{\mathrm{X}}^{\top}\varepsilon\left(\bm{T}_{\mathrm{W}}\bm{w}\right)+\bm{w}^{\top}\bm{T}_{\mathrm{W}}^{\top}\varepsilon\left(\bm{T}_{\mathrm{X}}\bm{x}\right)+\varepsilon\left(\bm{T}_{\mathrm{W}}\bm{w}\right)^{\top}\varepsilon\left(\bm{T}_{\mathrm{X}}\bm{x}\right)\right)^{2}$
$=$ $\ \left(\mathbb{E}_{\bm{w},\varepsilon\left(\bm{T}_{\mathrm{W}}\bm{w}\right)}\left\|\bm{X}^{\prime\top}\bm{T}_{\mathrm{X}}^{\top}\varepsilon\left(\bm{T}_{\mathrm{W}}\bm{w}\right)\right\|^{2}\right)+\left(\mathbb{E}_{\bm{x},\varepsilon\left(\bm{T}_{\mathrm{X}}\bm{x}\right)}\left\|\bm{W}^{\prime\top}\bm{T}_{\mathrm{W}}^{\top}\varepsilon\left(\bm{T}_{\mathrm{X}}\bm{x}\right)\right\|^{2}\right)$
Problem reduction. We can compute the optimal $\bm{T}_{\mathrm{W}}$ and $\bm{T}_{\mathrm{X}}$ individually for each of the non-negative terms and verify if $\bm{T}_{\mathrm{W}}=\bm{T}_{\mathrm{X}}^{-\top}$. The two terms have similar forms and can be optimized in similar ways. We focus on finding the optimal $\bm{T}_{\mathrm{X}}$ for the second term. Define the $d$-dimensional random variable $\bm{y}=\bm{W}^{\prime\top}\bm{x}\sim\mathcal{D}_{\mathrm{y}}$. Define the transformation matrix $\bm{T}=\bm{T}_{\mathrm{X}}\bm{W}^{\prime-\top}$. Then, the second term becomes
@@
Therefore, the two-sided quantization problem is reduced to finding the optimal $\bm{T}\in\mathbb{R}^{d\times d}$ that minimizes the one-sided quantization loss $\mathbb{E}_{\bm{y},\varepsilon\left(\bm{T}\bm{y}\right)}\left\|\bm{T}^{-1}\varepsilon\left(\bm{T}\bm{y}\right)\right\|^{2}$ and compute $\bm{T}_{\mathrm{X}}=\bm{T}\bm{W}^{\prime\top}$.
Transformation parameterization. Let the unknown orthogonal matrices $\bm{U}^{\prime},\bm{R}\in\mathbb{R}^{d\times d}$ and the unknown diagonal matrix $\bm{S}^{\prime}\in\mathbb{R}^{d\times d}$ be the singular value decomposition (SVD) of $\bm{T}\bm{U}\bm{S}=\bm{U}^{\prime}\bm{S}^{\prime}\bm{R}^{\top}$; thus, $\bm{T}$ can be parameterized as $\bm{T}=\bm{U}^{\prime}\bm{S}^{\prime}\bm{R}^{\top}\bm{S}^{-1}\bm{U}^{\top}$. Denote $\bm{S}=\operatorname{diag}\left(s_{1},\dots,s_{d}\right)$ and $\bm{S}^{\prime}=\operatorname{diag}\left(s^{\prime}_{1},\dots,s^{\prime}_{d}\right)$. Without loss of generality, assume $s_{1}\geq\dots\geq s_{d}>0$ and $s^{\prime}_{1}\geq\dots\geq s^{\prime}_{d}>0$.
@@
The optimal configuration of $\bm{T}$ for floating-point (FP) data types is $\bm{U}^{\prime}=\bm{H}$, $\bm{S}^{\prime}=\bm{S}^{\frac{1}{2}}$, and $\bm{R}=\mathbf{I}$. For integer (INT), the same configuration is optimal up to a $d^{o\left(1\right)}$ factor for zero-mean Gaussian/Laplacian distributed data and within a $d$ factor for any distribution.
Proof We provide detailed proofs for FP and INT in Sections˜2.3 and 2.2, respectively. For each data type, we apply a two-step proof strategy. We first provide a smooth modeling of the quantizer and treat the type casting errors as random noise. Secondly, we calculate the expected quantization error with respect to the parameterized transform and minimize this expectation by solving the corresponding optimization problem.
Some useful notations and basic results are as follows:
  • The second moment of $\bm{y}$ is $\mathbb{E}_{\bm{y}}\bm{y}\bm{y}^{\top}=\bm{W}^{\prime\top}\left(\mathbb{E}_{\bm{x}}\bm{x}\bm{x}^{\top}\right)\bm{W}^{\prime}=\bm{W}^{\prime\top}\bm{X}^{\prime}\bm{X}^{\prime\top}\bm{W}^{\prime}=\bm{U}\bm{S}^{2}\bm{U}^{\top}$.
  • The second moment of $\bm{T}\bm{y}$ is $\mathbb{E}_{\bm{y}}\bm{T}\bm{y}\left(\bm{T}\bm{y}\right)^{\top}=\bm{T}\left(\mathbb{E}_{\bm{y}}\bm{y}\bm{y}^{\top}\right)\bm{T}^{\top}=\bm{U}^{\prime}\bm{S}^{\prime 2}\bm{U}^{\prime\top}$.
  • The expected squared norm of $\bm{y}$ is $\mathbb{E}_{\bm{y}}\left\|\bm{y}\right\|^{2}=\operatorname{tr}\left(\mathbb{E}_{\bm{y}}\bm{y}\bm{y}^{\top}\right)=\operatorname{tr}\left(\bm{U}\bm{S}^{2}\bm{U}^{\top}\right)=\operatorname{tr}\left(\bm{S}^{2}\right)$.
  • Similarly, the expected squared norm of $\bm{T}\bm{y}$ is $\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|^{2}=\operatorname{tr}\left(\bm{S}^{\prime 2}\right)$.
  • Because $\bm{S}\succ\bm{0}$, $\left(\operatorname{tr}\left(\bm{S}\right)\right)^{2}\geq\operatorname{tr}\left(\bm{S}^{2}\right)$. By Jensen’s inequality, $\operatorname{tr}\left(\bm{S}^{2}\right)\geq d^{-1}\left(\operatorname{tr}\left(\bm{S}\right)\right)^{2}$.
@@ 2.2.1 Quantization Error Modeling
Theorem 2 (FP Quantization Error Modeling)
For a vector $\bm{\alpha}\in\mathbb{R}^{d}$, we can model the quantization error $\varepsilon\left(\bm{\alpha}\right)=\operatorname{diag}\left(\bm{\eta}\right)\bm{\alpha}$ with $\bm{\eta}=\left[\eta_{1},\dots,\eta_{d}\right]^{\top}\in\mathbb{R}^{d}$ being a random vector of i.i.d. samples $\eta\sim\mathcal{D}_{\mathrm{\eta}}$ and $\mathbb{E}_{\eta}\eta=0$.
Proof The FP format $\mathrm{E}a\mathrm{M}b$ is defined as a tuple $\left(s,e,m\right)$ with a sign $s\in\left\{0,1\right\}$, an exponent $e\in\left\{0,\dots,2^{a}-1\right\}$, and a mantissa $m\in\left\{0,\dots,2^{b}-1\right\}$. Assume we do not need to consider the subnormal and special number cases. The tuple $\left(s,e,m\right)$ represents the number $x_{\mathrm{E}a\mathrm{M}b}=\left(-1\right)^{s}2^{e-2^{a-1}+1}\left(1+2^{-b}m\right)$.
Using the approximation $1+\left(\cdot\right)\approx 2^{\left(\cdot\right)}$ for $2^{-b}m\in\left[0,1\right]$, we define a smoothed version of $\mathrm{E}a\mathrm{M}b$, named $\mathrm{SE}a\mathrm{M}b$, as $x_{\mathrm{SE}a\mathrm{M}b}=\left(-1\right)^{s}2^{e-2^{a-1}+1}2^{2^{-b}m}=\left(-1\right)^{s}2^{2^{-b}\left(2^{b}e+m\right)-2^{a-1}+1}$ with $2^{b}e+m=2^{b}\left(\log_{2}\left|x_{\mathrm{SE}a\mathrm{M}b}\right|+2^{a-1}-1\right)$ being a $a+b$ bit unsigned integer (concatenating the binary representations of $e$ and $m$). The minimum of $\left|x_{\mathrm{SE}a\mathrm{M}b}\right|$ is $\min\left|x_{\mathrm{SE}a\mathrm{M}b}\right|=2^{-2^{a-1}+1}$ and the maximum is $\max\left|x_{\mathrm{SE}a\mathrm{M}b}\right|=2^{2^{-b}\left(2^{a+b}-1\right)-2^{a-1}+1}=2^{2^{a-1}-2^{-b}+1}$.
Denote $\mathrm{SE}a\mathrm{M}b\left(x\right)$ as the quantizer that casts $x\in\left[-\max\left|x_{\mathrm{SE}a\mathrm{M}b}\right|,\max\left|x_{\mathrm{SE}a\mathrm{M}b}\right|\right]$ to the nearest number in the $\mathrm{SE}a\mathrm{M}b$ type. Assume $e\in\left(-\infty,2^{a}-1\right]$ so that the smaller values can be represented more precisely. $\mathrm{SE}a\mathrm{M}b\left(x\right)$ can be approximated using integer rounding in the logarithm space and first-order expansion. For $x\neq 0$,
$\mathrm{SE}a\mathrm{M}b\left(x\right)\approx$ $\ \mathrm{sgn}\left(x\right)2^{2^{-b}\left\lfloor 2^{b}\left(\log_{2}\left|x\right|+2^{a-1}-1\right)\right\rceil-2^{a-1}+1}=\mathrm{sgn}\left(x\right)2^{2^{-b}\left\lfloor 2^{b}\log_{2}\left|x\right|\right\rceil}$ (9)
For $x=0$, $\mathrm{SE}a\mathrm{M}b\left(0\right)\approx\lim_{x\to 0}x\left(1+\left(\left\lfloor 2^{b}\log_{2}\left|x\right|\right\rceil-2^{b}\log_{2}\left|x\right|\right)2^{-b}\ln 2\right)=0$. We model the casting error $\mathrm{SE}a\mathrm{M}b\left(x\right)-x=\left(\ln 2\right)2^{-b}\xi x$ where $x\in\left[-\max\left|x_{\mathrm{SE}a\mathrm{M}b}\right|,\max\left|x_{\mathrm{SE}a\mathrm{M}b}\right|\right]$ and the random variable $\xi\sim\mathrm{Uniform}\left(-2^{-1},2^{-1}\right)$.

### v2

@@ 4 Theoretical Derivation
In this section, we focus on the question of finding optimal transforms that minimize the loss of one block $\ell_{\left(i\right)}$ in Equation˜5. To build intuition for the proof, Figure˜1 visualizes how different transforms reshape the expected quantization error under FP and INT block quantizers, in a 2D toy setting. Throughout, we omit the block subscript $\left(i\right)$ to simplify notation. Also, the detailed derivation steps of some equations are deferred to Appendix Section˜A.3.
Figure 1: 2D illustration of how transforms shape the one-sided quantization error $\|\bm{T}^{-1}\varepsilon\left(\bm{T}\bm{y}\right)\|^{2}$ in Equation˜14 under FP and INT block quantizers. Each row corresponds to a different transform $\bm{T}$ (identity; Hadamard: 45° rotation; calibrated Hadamard: equally spreading energy into each dimension; whitening; WUS: WUSH without Hadamard; WUSH), and columns show FP and INT quantization in both original and transformed space. The dotted grid shows the basis induced by the transform. Two example points are shown in detail with their induced quantization noise support (Sections˜4.2.1 and 4.3.1) represented as shaded parallelograms, and the corresponding average (expected) error magnitude (Equation˜14) represented as the radii of circles (not necessarily proportional to the area of parallelograms). For each point on the plane, we visualize its expected error change (compared to that of the identity transform) calculated in the original space, with blue/red background heatmaps indicating lower/higher and dashed lines indicating the equilibrium set where the expected error is unchanged. No transform can reduce the error for all points on the plane, otherwise we would be able to repeatedly apply that transform until the error reaches zero. Thus, the key idea is to reduce the error around typical data distributions. The black ellipse depicts a representative data distribution contour. Overall, WUSH more effectively reduces the error by aligning the ellipse’s major axis to the error reduction (blue) regions.
@@ 4.1 Problem Setup and General Proof Approach
Probabilistic reformulation. We reformulate the problem from a probabilistic perspective. We split the matrices $\bm{W}=\left[\bm{w}_{1},\dots,\bm{w}_{d_{\mathrm{out}}}\right]$ and $\bm{X}=\left[\bm{x}_{1},\dots,\bm{x}_{d_{\mathrm{batch}}}\right]$ into columns $\bm{w}_{k},\bm{x}_{k}\in\mathbb{R}^{d}$ and treat the columns as i.i.d. samples from $d$-dimensional independent distributions $\bm{w}\sim\mathcal{D}_{\mathrm{w}}$ and $\bm{x}\sim\mathcal{D}_{\mathrm{x}}$, respectively. We motivate this multivariate modeling choice in Appendix Section˜A.2. We can reinterpret the blockwise loss in Equation˜5 as
$\ell=\mathbb{E}_{\bm{w},\bm{x}}\left(q\left(\bm{T}_{\mathrm{W}}\bm{w}\right)^{\top}q\left(\bm{T}_{\mathrm{X}}\bm{x}\right)-\bm{w}^{\top}\bm{x}\right)^{2}.$ (11)
Similarly, we can also reinterpret  Equation˜6 so that the matrices $\bm{W}^{\prime}$ and $\bm{X}^{\prime}$ satisfy
$\bm{W}^{\prime}\bm{W}^{\prime\top}$ $\ =\mathbb{E}_{\bm{w}}\bm{w}\bm{w}^{\top},$ (12)
$\bm{X}^{\prime}\bm{X}^{\prime\top}$ $\ =\mathbb{E}_{\bm{x}}\bm{x}\bm{x}^{\top},$
respectively.
Unbiased quantization. We assume that $q$ is a stochastic and unbiased quantizer, i.e., for a vector $\bm{\alpha}\in\mathbb{R}^{d}$, the quantization error $\varepsilon\left(\bm{\alpha}\right)=q\left(\bm{\alpha}\right)-\bm{\alpha}$ is a random vector with $\mathbb{E}_{\varepsilon\left(\bm{\alpha}\right)}\varepsilon\left(\bm{\alpha}\right)=\bm{0}$. We further constrain $\bm{T}_{\mathrm{W}}=\bm{T}_{\mathrm{X}}^{-\top}$ such that $q\left(\bm{T}_{\mathrm{W}}\bm{w}\right)^{\top}q\left(\bm{T}_{\mathrm{X}}\bm{x}\right)$ is unbiased with respect to $\bm{w}^{\top}\bm{x}$. Then, using the first-order approximation, we can split $\ell$ in  Equation˜11 into two non-negative terms:
$\ell=\left(\mathbb{E}_{\bm{w},\varepsilon\left(\bm{T}_{\mathrm{W}}\bm{w}\right)}\left\|\bm{X}^{\prime\top}\bm{T}_{\mathrm{W}}^{-1}\varepsilon\left(\bm{T}_{\mathrm{W}}\bm{w}\right)\right\|^{2}\right)+\left(\mathbb{E}_{\bm{x},\varepsilon\left(\bm{T}_{\mathrm{X}}\bm{x}\right)}\left\|\bm{W}^{\prime\top}\bm{T}_{\mathrm{X}}^{-1}\varepsilon\left(\bm{T}_{\mathrm{X}}\bm{x}\right)\right\|^{2}\right).$ (13)
The detailed derivation is given in Appendix Equation˜37.
Problem reduction. We can compute the optimal $\bm{T}_{\mathrm{W}}$ and $\bm{T}_{\mathrm{X}}$ individually for each of the non-negative terms and verify that $\bm{T}_{\mathrm{W}}=\bm{T}_{\mathrm{X}}^{-\top}$. The two terms have similar forms and can be optimized in similar ways. We focus on finding the optimal $\bm{T}_{\mathrm{X}}$ for the second term. Define the $d$-dimensional random variable $\bm{y}=\bm{W}^{\prime\top}\bm{x}$ and denote its distribution as $\bm{y}\sim\mathcal{D}_{\mathrm{y}}$. Define the transformation matrix $\bm{T}=\bm{T}_{\mathrm{X}}\bm{W}^{\prime-\top}$. Then, the second term in Equation˜13 becomes
@@
Therefore, the two-sided quantization problem in Equation˜11 is reduced to finding the optimal $\bm{T}\in\mathbb{R}^{d\times d}$ that minimizes the one-sided quantization loss $\mathbb{E}_{\bm{y},\varepsilon\left(\bm{T}\bm{y}\right)}\left\|\bm{T}^{-1}\varepsilon\left(\bm{T}\bm{y}\right)\right\|^{2}$ and compute
$\bm{T}_{\mathrm{X}}=\bm{T}\bm{W}^{\prime\top}.$ (15)
Transformation parameterization. Let the unknown orthogonal matrices $\bm{U}^{\prime},\bm{R}\in\mathbb{R}^{d\times d}$ and the unknown diagonal matrix $\bm{S}^{\prime}\in\mathbb{R}^{d\times d}$ be the singular value decomposition (SVD) of $\bm{T}\bm{U}\bm{S}=\bm{U}^{\prime}\bm{S}^{\prime}\bm{R}^{\top}$ with $\bm{U},\bm{S}$ defined in Equation˜7; thus, $\bm{T}$ can be parameterized as
$\bm{T}=\bm{U}^{\prime}\bm{S}^{\prime}\bm{R}^{\top}\bm{S}^{-1}\bm{U}^{\top}.$ (16)
Denote $\bm{S}=\operatorname{diag}\left(s_{1},\dots,s_{d}\right)$ and $\bm{S}^{\prime}=\operatorname{diag}\left(s^{\prime}_{1},\dots,s^{\prime}_{d}\right)$. Without loss of generality, assume $s_{1}\geq\dots\geq s_{d}>0$ and $s^{\prime}_{1}\geq\dots\geq s^{\prime}_{d}>0$. A useful property of $\bm{S}$ is
$\operatorname{tr}\left(\bm{S}^{2}\right)\geq d^{-1}\left(\operatorname{tr}\left(\bm{S}\right)\right)^{2}\geq d^{-1}\operatorname{tr}\left(\bm{S}^{2}\right)$ (17)
with detailed steps in Equation˜38. The equality $\operatorname{tr}\left(\bm{S}^{2}\right)=d^{-1}\left(\operatorname{tr}\left(\bm{S}\right)\right)^{2}$ is attained when all $s_{i}$ are the same (all singular values are inliers). The equality $d^{-1}\left(\operatorname{tr}\left(\bm{S}\right)\right)^{2}=d^{-1}\operatorname{tr}\left(\bm{S}^{2}\right)$ is attained when $\operatorname{tr}\left(\bm{S}\right)\to s_{1}$ (a singular value is an extreme outlier).
@@
The optimal configuration of $\bm{T}$ for floating-point (FP) data types is $\bm{U}^{\prime}=\bm{H}$, $\bm{S}^{\prime}=\bm{S}^{\frac{1}{2}}$, and $\bm{R}=\mathbf{I}$. For integer (INT) types, the same configuration is optimal up to a $d^{o\left(1\right)}$ factor for zero-mean multivariate Gaussian/Laplacian distributed data and within a $d$ factor for any distribution.
Proof We provide detailed proofs for FP and INT in Sections˜4.2 and 4.3, respectively. For each data type, we apply a two-step proof strategy. We first (Sections˜4.2.1 and 4.3.1) provide a smooth modeling of the quantizer and treat the type casting errors as random noise. Secondly (Sections˜4.2.2 and 4.3.2), we calculate the expected quantization error, Equation˜14, with respect to the parameterized transform, Equation˜16, and minimize this expectation by solving the corresponding optimization problem.
The optimal $\bm{T}_{\mathrm{X}}$, aka the WUSH construction $\bm{T}_{\mathrm{wush}}$ in Equation˜8, is obtained by applying Theorems˜1, 15 and 16.
@@ 4.2.1 Quantization Error Modeling
For a number in FP types, the quantization error tends to be proportional to its absolute value. Formally, for a vector $\bm{\alpha}\in\mathbb{R}^{d}$, we can model the quantization error as
$\varepsilon\left(\bm{\alpha}\right)=\operatorname{diag}\left(\bm{\eta}\right)\bm{\alpha}$ (18)
where $\bm{\eta}=\left[\eta_{1},\dots,\eta_{d}\right]^{\top}\in\mathbb{R}^{d}$ is a random vector of i.i.d. samples from the distribution $\eta\sim\mathcal{D}_{\mathrm{\eta}}$ with $\mathbb{E}_{\eta}\eta=0$. We provide more rigorous justifications for why this modeling makes sense in Appendix Section˜A.4.2.
4.2.2 Loss Minimization
Using the quantization error modeling in Equation˜18, the minimization objective in Equation˜14 becomes
$\ \mathbb{E}_{\bm{y},\varepsilon\left(\bm{T}\bm{y}\right)}\left\|\bm{T}^{-1}\varepsilon\left(\bm{T}\bm{y}\right)\right\|^{2}=\left(\mathbb{E}_{\eta}\eta^{2}\right)\operatorname{tr}\left(\bm{T}^{-1}\left(\left(\bm{U}^{\prime}\bm{S}^{\prime 2}\bm{U}^{\prime\top}\right)\odot\mathbf{I}\right)\bm{T}^{-\top}\right)$ (19)
with $\odot$ representing elementwise multiplication, and with detailed steps in Equation˜39. The lower bound of the trace term in Equation˜19 is
$\operatorname{tr}\left(\bm{T}^{-1}\left(\left(\bm{U}^{\prime}\bm{S}^{\prime 2}\bm{U}^{\prime\top}\right)\odot\mathbf{I}\right)\bm{T}^{-\top}\right)\geq d^{-1}\left(\operatorname{tr}\left(\bm{S}\right)\right)^{2}$ (20)
and the equality can be attained by choosing $\bm{U}^{\prime}=\bm{H}$, $\bm{S}^{\prime}=\bm{S}^{\frac{1}{2}}$, and $\bm{R}=\mathbf{I}$, with detailed steps in Equations˜40 and 41.
4.2.3 Discussion
Under our smooth modeling in Equation˜18, for any orthogonal $\bm{T}$ (including the identity $\bm{T}=\mathbf{I}$ and the Hadamard $\bm{T}=\bm{H}$), the minimization objective becomes trivially
$\mathbb{E}_{\bm{y},\varepsilon\left(\bm{T}\bm{y}\right)}\left\|\bm{T}^{-1}\varepsilon\left(\bm{T}\bm{y}\right)\right\|^{2}=\left(\mathbb{E}_{\eta}\eta^{2}\right)\operatorname{tr}\left(\bm{S}^{2}\right)$ (21)
with detailed steps in Equation˜42. Thus, orthogonal transforms will not be helpful for reducing the quantization error. Comparing the trace terms in Equations˜20 and 21 using the inequality Equation˜17, the WUSH transform can reduce the error by at most $d$ times in the extreme outlier scenario. Also note that the non-trivial choices $\bm{U}^{\prime}=\bm{H}$ and $\bm{S}^{\prime}=\bm{S}^{\frac{1}{2}}$ are both essential for reducing the error. Either trivial choices of $\bm{U}^{\prime}=\mathbf{I}$ or $\bm{S}^{\prime}=\mathbf{I}$ will lead to the same suboptimal trace term
$\operatorname{tr}\left(\bm{T}^{-1}\left(\left(\bm{U}^{\prime}\bm{S}^{\prime 2}\bm{U}^{\prime\top}\right)\odot\mathbf{I}\right)\bm{T}^{-\top}\right)=\operatorname{tr}\left(\bm{S}^{2}\right)$ (22)
as that in the case of the orthogonal $\bm{T}$ in Equation˜21, with detailed steps in Equations˜43 and 44. Figure˜1 (left two columns) also visualizes these results.
4.3 Proof for Integer (INT) Data Types
4.3.1 Quantization Error Modeling
For a number in INT types, the quantization error tends to be proportional to the maximum absolute value within a quantization group. Formally, for a vector $\bm{\alpha}\in\mathbb{R}^{d}$, we can model the quantization error as
$\varepsilon\left(\bm{\alpha}\right)=\left\|\bm{\alpha}\right\|_{\infty}\bm{\eta}$ (23)
where $\bm{\eta}=\left[\eta_{1},\dots,\eta_{d}\right]^{\top}\in\mathbb{R}^{d}$ is a random vector of i.i.d. samples from the distribution $\eta\sim\mathcal{D}_{\mathrm{\eta}}$ with $\mathbb{E}_{\eta}\eta=0$. We provide more rigorous justifications for why this modeling makes sense in Appendix Section˜A.4.1.
4.3.2 Loss Minimization
Using the quantization error modeling in Equation˜23, the minimization objective in Equation˜14 becomes
$\mathbb{E}_{\bm{y},\varepsilon\left(\bm{T}\bm{y}\right)}\left\|\bm{T}^{-1}\varepsilon\left(\bm{T}\bm{y}\right)\right\|^{2}=\left(\mathbb{E}_{\eta}\eta^{2}\right)\left\|\bm{T}^{-1}\right\|_{\mathrm{F}}^{2}\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|_{\infty}^{2}$ (24)
with detailed steps in Equation˜45.
We can obtain a lower bound for the term $\left\|\bm{T}^{-1}\right\|_{\mathrm{F}}^{2}$ in Equation˜24.
$\left\|\bm{T}^{-1}\right\|_{\mathrm{F}}^{2}\geq\operatorname{tr}\left(\bm{S}^{2}\bm{S}^{\prime-2}\right),$ (25)
and the equality is attained when $\bm{R}=\mathbf{I}$, with detailed steps in Equations˜46 and 47.
Next, we obtain near-optimal lower and upper bounds for the term $\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|^{2}_{\infty}$ in Equation˜24 by choosing $\bm{U}^{\prime}=\bm{H}$.